OpenStudy (anonymous):

Express The square root of negative 72. in i notation.

OpenStudy (anonymous):

\[\sqrt{-72}\]

OpenStudy (anonymous):

i bet you know this

OpenStudy (solomonzelman):

i72, this is too EZ! and you know the answer to this question? ..... Do you teach math anywhere?

OpenStudy (anonymous):

actually if it was me, i would write \[6\sqrt{2}i\] since \(\sqrt{72}=6\sqrt2\)

OpenStudy (solomonzelman):

OK, I agree!

OpenStudy (anonymous):

Im trying still i did this \[8i \pm \sqrt{3}\]

OpenStudy (anonymous):

there is no sum in this

OpenStudy (anonymous):

Ohh ok

OpenStudy (anonymous):

i forgot how to do that sqrt thing

OpenStudy (anonymous):

\[6 \sqrt{2}\]

OpenStudy (anonymous):

\(72=36\times 2\)

OpenStudy (anonymous):

Ohh and i take sqrt of 36 and i get 6 oh ok thanks again

OpenStudy (anonymous):

Oh and Mr. satellite when i simplify \[i ^{18}\]

OpenStudy (anonymous):

Is it \[i ^{16}\times i ^{2}?\]

OpenStudy (anonymous):

i guess so i would ask the following simpler question what is the integer remainder when you divide 18 by 4?

OpenStudy (anonymous):

\[-i?\]

OpenStudy (anonymous):

let me ask again forget complex numbers for a second what is the remainder when you divide 18 by 4?

OpenStudy (anonymous):

or -1 i think -1

OpenStudy (anonymous):

its 4.5

OpenStudy (anonymous):

once more what is the "remainder" when you divide 18 by 4?

OpenStudy (anonymous):

ohh man i fell dumb its 2

OpenStudy (anonymous):

whew!

OpenStudy (anonymous):

then \[i^{18}=i^2=-1\]

OpenStudy (anonymous):

sorry my head is just acting this stuff is hard to me

OpenStudy (anonymous):

similarly \(i^{203}=i^3=-i\)

OpenStudy (anonymous):

\(i^{12345}=i^1=i\) and so on

OpenStudy (anonymous):

i was looking at that and it seems like odd nubers will be i and even numbers will be 1

OpenStudy (anonymous):

not numbers the exponents

OpenStudy (anonymous):

but then i just need to find weather its positive or negative

OpenStudy (anonymous):

take the integer remainder when you divide the exponent by 4, that is all

OpenStudy (anonymous):

\[i^0=1,i^1=i, i^2=-1,i^3=-i\] it has to be one of those

OpenStudy (anonymous):

if the remainder is 3, it is \(-i\) if the remainder is 2, it is \(-1\) if the remainder is 1, it is \(i\) and if the remainder is 0, i.e. if the exponent is divisible by 4, then it is 1

OpenStudy (anonymous):

Ok thanks im worn out now i need sleep gn